The electron drift arises due to the force experienced by electrons in the electric field inside the conductor.

But force should cause acceleration.

Why then do the electrons aquire a steady average drift speed?

1
Expert's answer
2021-12-14T10:33:41-0500

(1)Each free electron does accelerate increasing its drift speed after collision but starts to accelerate and increases its drift speed

(2)again only to suffer a collision again and so on on the average, thereofre electrons acquire only a drift speed.


We know that

"F=ma 
\\F=qE"

Both equation equal


"ma=qE"

"a=\frac{qE}{m}"

"q=-e"

"a=\frac{-eE}{m}"

Drinft velocity

"J=neAv_d"

"V_d=\frac{J}{neA}"